Mike200409:

The equation y = negative StartFraction 1 Over 20 EndFraction x + 10 represents the gallons of gasoline that remain in Michelle’s car after she drives x miles. Which graph represents this equation?

itsme:

Given : The equation y = -1/20 x + 10 . To find : Which graph represents this equation. Solution : We have given an equation y = -1/20 x + 10 .We can see it is a slope intercept form here slope = 1/20. Here, y - intercept = 10 ( Graph line cut at 10 on y axis )

supie:

y = negative StartFraction 1 Over 20 EndFractionx + 10 represents\(\LARGE\longrightarrow\) \(\LARGE y=-\frac{1}{20}x+10\) ?

lilianamendez2:

it say y = negative StartFraction 1 Over 20 EndFraction x + 10 represent the gallons of gasoline that remain in Michelle’s car after she drives x miles.

lilianamendez2:

and it say Which graph represents this equation? but they dont put choices